3. The brakes of an automobile traveling with a velocity of 50 ft/s are suddenly applied. If the automobile comes to a stop after 5 s, what is its acceleration?


1
Expert's answer
2021-09-14T09:36:22-0400

By definition, the acceleration is given as a rate of speed change:


"a = \dfrac{v_f-v_i}{t}"

where "v_f = 0ft\/s" is the finall speed (car stops), "v_i = 50ft\/s" is initial speed, and "t = 5s" is the time. Thus, obtain:


"a = \dfrac{0-50ft/s}{5s} = -10ft/s^2"

Answer. "-10ft\/s^2".